# Identifying User Goals from UI Trajectories

### v2

@@
Abstract
Autonomous agents that interact with graphical user interfaces (GUIs) hold significant potential for enhancing user experiences. To further improve these experiences, agents need to be personalized and proactive. By effectively comprehending user intentions through their actions and interactions with GUIs, agents will be better positioned to achieve these goals. This paper introduces the task of goal identification from observed UI trajectories, aiming to infer the user’s intended task based on their GUI interactions. We propose a novel evaluation metric to assess whether two task descriptions are paraphrases within a specific UI environment. By Leveraging the inverse relation with the UI automation task, we utilized the Android-In-The-Wild and Mind2Web datasets for our experiments. Using our metric and these datasets, we conducted several experiments comparing the performance of humans and state-of-the-art models, specifically GPT-4 and Gemini-1.5 Pro. Our results show that Gemini performs better than GPT but still underperforms compared to humans, indicating significant room for improvement.
Identifying User Goals from UI Trajectories
Omri Berkovitchthanks: Equal contribution, 1  Sapir Caduri11footnotemark: 1, 1  Noam Kahlon1  Anatoly Efros1 Avi Caciularu1Ido Dagan1, 2 1Google Research  2Bar-Ilan University {berkovitchomri,sapir,kahlonn,talef,avica,idodagan}@google.com
1 Introduction
Autonomous agents that interact with GUIs to complete tasks for users have drawn significant interest Hong et al. (2023); Gur et al. (2023); Yang et al. (2023b). These agents interpret natural language task descriptions provided by a user and iteratively interact with GUIs to complete the desired task. Enhancing an agent’s ability to understand user actions and infer goals within the GUI environment has the potential to significantly increase its utility for users by providing more personalized and effective assistance Li et al. (2024).
@@ 1 Introduction
Consider the scenario in Figure 1 where a user books flight tickets online for a vacation. An ideal agent would first observe and understand these actions, then proactively suggest booking a hotel for the same dates, and make the vacation dates visible in the calendar. This kind of assistance requires the agent to understand the underlying goals driving user interactions. Additionally, by identifying user goals, we can augment the agent’s memory, making it more personal and useful for future tasks. We frame this challenge as the inverse of the known UI automation task: identifying the hidden user goal from its observed UI interactions Li et al. (2023); Wen et al. (2023). While intent identification has been explored in other domains Hu et al. (2009); Gupta et al. (2014), our work is the first, to the best of our knowledge, to focus specifically on identifying user goals from UI interactions.
@@
Figure 1: An example of a user performing a flight booking task. The agent first observes the UI interactions, comprehends the task’s essence, and then offers help with related tasks, like booking a hotel and blocking calendar dates. We focus on the first part, comprehending the task by observing the UI interactions.
Identifying user goals from UI interactions is challenging. Although ensuring a correct goal is achievable through the observed actions is essential, this alone is not enough. The fact that multiple tasks can often lead to the same actions making the task ambiguous, further highlighting the complexity of the task and the need for robust methods. Additionally, the involvement of multiple modalities, such as visual and text, increases complexity.
In this paper, we explore the task of goal identification from observed UI trajectories. In particular, we introduce a robust manual and automatic evaluation protocol. Our proposed metric primarily assesses whether two task descriptions are paraphrases within the UI context while also enabling us to quantify the ambiguity. We conducted extensive experiments in both web and Android UIs, leveraging existing UI automation datasets by swapping the roles of input and output. We compared the performance of humans and state-of-the-art multi-modal models on this task and found that the models significantly underperform, indicating substantial room for improvement. Common errors and potential mitigations are highlighted, paving the way for future research.
Overall our main contributions are as follows: (1) presenting the first study on the task of goal identification from UI trajectories, including task formulation and definitions; (2) introducing a novel manual and automatic evaluation methodology and (3) assessing the performance of both humans and models on this task, providing a comprehensive analysis of our findings.
2 Task Definition
@@
Given an observed UI trajectory performed by a user with the intention to complete a certain task, our goal is to recover the user’s original intent from the observable trajectory. As mentioned in § 1, this intent identification setting is effectively the inverse problem of the known UI Automation task. We therefore adopt their input and output definitions, swapping their roles, which enables the use of UI automation datasets for the intent identification task as well. Accordingly, our input is a UI trajectory - a sequence of steps performed by the user. Each step consists of a snapshot of the UI environment at that moment, along with the corresponding action the user took at that step. From this trajectory, our goal is to generate a natural language description that accurately captures the user’s intended task. Within the scope of this paper, we address the core setting of the intent identification task and therefore assume that the observed UI trajectory indeed successfully fulfills the underlying user intent.
@@
The intent identification task, similar to other text generation tasks like summarization where multiple valid outputs can exist, is inherently ambiguous, mostly because the same trajectory may fulfill multiple intents. For example, when observing a Sushi restaurant booking, the user might have asked for that specific Sushi place, or more broadly for some restaurant in that area, or of that cuisine type. When a model identifies intent from trajectory, we expect it to predict the most likely one.
3 Evaluation Methodology
@@
This section outlines our evaluation methodology. Given an input UI trajectory and a corresponding gold task description, we assess whether a predicted task description matches the gold reference. We start with necessary definitions (Section 3.1), followed by a human evaluation protocol (Section 3.2) and a language model-based automatic evaluation (Section 3.3).
3.1 Definitions
@@
Inspired by the taxonomy proposed by Zhou et al. (2023), we differentiate between fulfillment for information-seeking intents and transactional intents (e.g. purchasing an item, booking a flight or changing settings). For the latter, fulfillment is achieved upon successfully completing the specific requirement outlined in the task. For information-seeking intents, fulfillment is achieved when the trajectory provides the necessary information sought in the user intent. We note that a fulfilling trajectory may provide some additional information beyond the intent, if such additional information is inherently bundled in the UI environment together with the sought information (e.g. some UI environments may display review scores along with other related information).
@@
We next aim to specify a matching criterion between two task descriptions, specifically a predicted one and the corresponding gold-reference. Given two task descriptions A and B and a UI environment, we say that A satisfies B in that environment if every reasonable user trajectory that fulfills A would also fulfill B. In essence, this means that completing task A necessarily results in completing task B, making B a more general task than A in that UI environment. For instance, the task “Purchase the earliest train ticket to Edinburgh” satisfies the task “Purchase a ticket to Edinburgh” but not vice versa.
@@
3.2 Human Evaluation Protocol
As with many text generation tasks, human evaluation is essential due to the limited reliability of automatic evaluation metrics. In our case, the human evaluator observes the gold and predicted task descriptions, along with the corresponding trajectory, and assesses (1) whether the trajectory fulfills the predicted task, and (2) whether each of the predicted and gold task descriptions satisfies the other.
To verify the quality of our suggested metric, we sampled randomly 50 instances from Mind2Web dataset (see § 4.1) and generated task descriptions using two baseline models (see § 4.2) evaluated in our experiments. We measured the pairwise inter-annotator agreement between three of the authors yielding average Cohen’s Kappa of 0.79 for fulfillment judgement and 0.77 for satisfaction judgement. These agreement levels are in the higher range of the “Substantial agreement” level, thus we proceeded to manually evaluate each baseline model in our experiments § 5 by a single annotator.
3.3 Automatic Evaluation Metric
We propose utilizing a Large Multimodal Model (LMM) as an automatic evaluator for the satisfaction criteria. Recent advancements in LMMs, such as those demonstrated in He et al. (2024) and Pan et al. (2024), have shown promising results in employing GPT-4 Achiam et al. (2023) to assess task completion by autonomous agents. Building on this, we leverage the latest GPT-4o model as the automatic evaluator to determine whether two task descriptions, are mutually satisfied in the context of the trajectory. The specific prompt used is detailed in Appendix D. Our analysis indicates that the proposed automatic evaluation generally agrees with the human annotators, with an F1 score of 0.75, and can be a sufficient substitute for manual evaluation, enabling quick development cycles.
Dataset Non Partial  Match ($\uparrow$)
 Match  Match
@@ 3.3 Automatic Evaluation Metric
AitW H   vs. G 0.12 0.32 0.56
  H1 vs. H2 0.05 0.20 0.75
Table 1: Manual evaluation of human generated task descriptions, at the different levels of match criteria (determined by the satisfaction relation) between annotators (Hi) and gold task descriptions (G), as well as among annotators. (H) represents average scores.
4 Datasets and Baseline Models
@@
Dataset Model Fulfillment Non Match Partial Match  Match ($\uparrow$) Non Match Partial Match  Match ($\uparrow$)
@@
AitW GPT-4 0.65 0.50 0.20 0.30 0.62 0.19 0.19
Gemini-1.5 0.89 0.19 0.31 0.50 0.36 0.29 0.35
Table 2: Manual and automatic evaluation scores of fulfillment relation and the different levels of match criteria (determined by the satisfaction relation), between model predictions and gold task descriptions.
4.1 Evaluated Datasets
Given that our task is the inverse of UI automation task, leveraging the datasets created for UI automation is a natural choice. In these datasets, humans perform actions within a UI environment, based on a given task description.
@@ 4.1 Evaluated Datasets
For our experiments, we explore two UI environments: web and Android. We utilize the Mind2Web dataset Deng et al. (2023) for the web environment, as it is the most widely used benchmark for autonomous web agents. The dataset was created by curating diverse tasks across popular websites, with annotators performing a series of actions to complete the goal. For Android, we used the Android in the Wild (AitW) Rawles et al. (2023) dataset, notable for its diverse task descriptions spanning various domains. Further details in Appendix A.
4.2 Models
@@
Given the multimodal nature of UI trajectories, our models must be adept at interpreting both text and images. We selected two state-of-the-art LMMs, Gemini 1.5 Pro Reid et al. (2024) and GPT-4 Achiam et al. (2023). These models are at the forefront of handling combined text and image inputs and offer a sufficiently large context window, which is necessary for our experiments.
@@
Our experiments focused on in-context learning, employing both zero-shot and few-shot prompting. The model was guided through a Chain-of-Thought Wei et al. (2022) process, to analyze the trajectory in a step-by-step manner (details in Appendix D).
@@
In web, inspired by SeeAct Zheng et al. (2024) we drew a red bounding box around the element that the user interacted with to guide the model’s attention, as well as to break ambiguity where textual action descriptions were not sufficient. Additionally, we truncated lengthy web page images based on the bounding box position to further focus the model’s attention (details in Appendix B). For Android, no special modifications were necessary, as actions were overlaid on the screenshots and the screenshots were limited to the visible area.
5 Experiments
5.1 Human Performance Evaluation

### v3

@@
Abstract.
Identifying underlying user goals and intents has been recognized as valuable in various personalization-oriented settings, such as personalized agents, improved search responses, advertising, user analytics, and more. In this paper, we propose a new task—goal identification from observed UI trajectories—aiming to infer the user’s detailed intentions when performing a task within UI environments. To support this task, we also introduce a novel evaluation methodology designed to assess whether two intent descriptions can be considered paraphrases within a specific UI environment. Furthermore, we demonstrate how this task can leverage datasets designed for the inverse problem of UI automation, utilizing Android and web datasets for our experiments. To benchmark this task, we compare the performance of humans and state-of-the-art models, specifically GPT-4 and Gemini-1.5 Pro, using our proposed metric. The results reveal that both Gemini and GPT underperform relative to human performance, underscoring the challenge of the proposed task and the significant room for improvement. This work highlights the importance of goal identification within UI trajectories, providing a foundation for further exploration and advancement in this area.
Intent Understanding, Proactive Agents, Personalized Agents, Multimodal Interaction
journalyear: 2025copyright: ccconference: Companion Proceedings of the ACM Web Conference 2025; April 28-May 2, 2025; Sydney, NSW, Australiabooktitle: Companion Proceedings of the ACM Web Conference 2025 (WWW Companion ’25), April 28-May 2, 2025, Sydney, NSW, Australiadoi: 10.1145/3701716.3717525isbn: 979-8-4007-1331-6/2025/04ccs: Human-centered computing User modelsccs: Human-centered computing User studiesccs: Computing methodologies Intelligent agentsccs: Human-centered computing Field studies
1. Introduction
Autonomous agents that interact with GUIs to complete tasks for users have drawn increasing interest (Hong et al., 2023; Gur et al., 2023; Yang et al., 2023a). These agents interpret user-provided instructions and iteratively interact with GUIs to complete the desired task. In this work, we propose empowering agents with improved personalization abilities, allowing them to identify the underlying goals of users from their observed activity within the GUI environment. Understanding the underlying goals and intents of users is a cornerstone of user modeling, enabling agents to proactively align their behavior with user preferences, habits, and long-term objectives. This capability marks a significant step toward achieving personalization, where agents can provide context-aware, effective assistance that meets individual user needs (Li et al., 2024).
@@ 1. Introduction
Consider the scenario in Figure 1 where a user books flight tickets for a vacation. An ideal agent would observe these actions, understand the underlying user goal, and then proactively suggest booking a hotel for the same dates, and make the dates visible in the calendar.
Our work extends a long line of research on recognizing user goals from their observed behaviour, including intent, activity, and plan recognition. However, most prior work addressed settings in which the input consists of natural language user utterances, like search queries, dialog utterances, or social media posts, while the task was perceived as structured classification, selecting a category label from a predefined list. Our work, to the best of our knowledge, is the first to identify user goals from UI interactions, while providing a natural language description of the user goal as output. We refer to Section 2 for an extensive overview.
In this paper, we first define the task of generating a natural language description of underlying user goals from observed UI trajectories, that is, from multi-modal (text and screen image) traces of user-system interactions. A challenging aspect of the task is its inherent ambiguity, since multiple user goals can often lead to the same UI activity. Second, we observe that our task can be framed as the inverse of the known UI automation task, where an agent needs to perform a sequence of UI actions given a natural language description of the user’s goal (Li et al., 2023; Wen et al., 2023). We further introduce manual and automatic evaluation protocols, assessing whether the predicted and gold task descriptions are paraphrases within the given UI context.
Subsequently, we conducted experiments over both web and Android UI sessions, leveraging existing UI automation datasets while swapping the input and output roles. Over these data, we compared and analyzed the performance of humans and state-of-the-art multi-modal models, showing that there is substantial room for modeling improvements in future work.
Overall, we offer the following contributions: (1) introducing and formalizing the task of goal identification from UI trajectories, highlighting its significance and potential applications; (2) demonstrating how existing datasets for UI automation can be adapted for this task by interpreting it as the ”reversed” process of the other; (3) proposing both manual and automatic evaluation methodologies for performance assessment; and (4) evaluating the performance of both humans and state-of-the-art models on this task.
We propose that these contributions would trigger research on this timely challenge.
@@
Figure 1. An example of a user performing a flight booking task. The agent first observes the UI interactions, comprehends the task’s essence, and then offers help with related tasks, like booking a hotel and blocking calendar dates. We focus on the first part, comprehending the task by observing the UI interactions.
[…]
The figure illustrates a sequence of UI interactions in a flight booking scenario. A user selects flight details like destination and dates, and the agent interprets these actions. Additional UI interactions for related tasks, such as booking a hotel and calendar updates, are shown. The agent observes and understands these UI trajectories.
2. Background and Related Work
In this section, we provide an overview of key related research areas that aim to learn personal user behaviors and preferences, specifically intent recognition, activity recognition, and plan recognition.
Intent Recognition
Intent recognition, also referred to as intent classification, is a well-established field focused on identifying user intentions based on specific inputs. Traditionally, most research in this domain has concentrated on textual inputs, such as individual messages or utterances, which are then categorized into predefined intent classes. For instance, studies like (Schuurmans and Frasincar, 2019; Kuchlous and Kadaba, 2020; Larson and Leach, 2022) have worked on improving intent classification within dialogue systems across various domains, focusing on the classification of short texts or single utterances into predetermined intent categories (e.g. “find a train from Barcelona to Madrid” would be classified into a system intent called ”find_train”). Others, classified social media posts to determine whether they express an intention to make a purchase (Gupta et al., 2014; Haque et al., 2019). To enhance understanding of user queries and infer its underlying intent, some works leverage external knowledge sources to get better results (Hu et al., 2009).
In addition to text-based inputs, multimodal methods have been developed to incorporate various types of inputs. For example, (Zhang et al., 2021) and (Gonzaga et al., 2021) investigate the use of both images and text: the former analyzes social news content to identify marketing intents and to classify intent topics, while the latter combines image and textual data to classify social media posts, aiming to identify the writers’ intent such as provocative, informative, promotive and more. Zhang et al. (2022) introduces a multimodal dataset for intent recognition in TV series, where inputs include visual, auditory, and textual data, while the outputs are classified into one of 25 possible intent categories.
Activity Recognition
Activity recognition, as defined by Sukthankar et al. (2014) and closely related to intent recognition, involves identifying specific human activities based on a series of observations, often through sensor data. While intent recognition focuses on understanding user intentions, activity recognition mostly concerned with classifying physical actions. Typically, systems in this domain are developed to classify these activities using sensory inputs. Khan et al. (2022) uses a neural network trained on 2D skeletal data captured by a motion sensor to classify human poses; Ijaz et al. (2022) integrated accelerometer signals along with skeletal data to recognize and categorize nurse activities into 12 distinct types. A prominent study (Kwapisz et al., 2011) developed a supervised learning algorithm that uses accelerometer signal from Android smartphones to classify physical activities like walking, jogging, sitting, and standing. Moreover, research such as (Liao et al., 2005) utilizes GPS traces to classify human activity and label significant locations, employing relational Markov networks to achieve this.
Plan Recognition
In a plan recognition problem, a system is given a series of actions performed by an agent and is expected to infer the overall plan or goal which explains those actions (Kautz et al., 1991). The distinction between activity recognition and plan recognition is the difference between recognizing a single activity and recognizing the relationships between a set of such activities that result in a complete plan or goal. In this field, foundational works like (Charniak and Goldman, 1993; Goldman et al., 2013) rely on structured inputs and a plan library — a collection of possible plans and their associated goals — to construct probabilistic models that infer the plan from observed actions. Alternative approaches, such as in (Hong, 2001), develop algorithms that do not depend on plan libraries; instead, they utilizing in-domain knowledge and applying this to UNIX system commands to understand broader user goal. A hybrid approach is demonstrated (Granada et al., 2017), where the authors combine CNN models to detect human activities from a video stream of meal preparation and apply symbolic reasoning with a plan library to recognize the overall plan (e.g., actions like ”breaking eggs” and ”mixing ingredients” might be recognized as part of a plan to create an omelet).
3. Task Definition
@@
Given an observed UI trajectory (see below) performed by a user with the intention to complete a certain task, our goal is to recover the user’s original intent from the observable trajectory. As mentioned in § 1, this setting is effectively the inverse problem of the known UI Automation task. We therefore adopt their input and output definitions, swapping their roles, which enables the use of UI automation datasets for our task as well.
Accordingly, our input is a UI trajectory — a sequence of individual interaction steps between the user and the system along the session. Each step consists of a snapshot of the UI content at that moment, along with the corresponding action the user took at that step (see Appendix A for the specific trajectory formats). From this trajectory, our goal is to generate a natural language description that accurately captures the user’s intended task. Within the scope of this paper, we address the core setting of the intent identification task and therefore assume that the observed UI trajectory indeed successfully fulfills the underlying user intent.
@@
The intent identification task, similar to other text generation tasks like summarization where multiple valid outputs can exist, is inherently ambiguous, mostly because the same trajectory may fulfill multiple intents, often due to varying specificity levels of the original user intent (as captured in the dataset). For example, when observing a Sushi restaurant booking, the user might have asked for that specific Sushi place, or more broadly for some restaurant in that area, or of that cuisine type. When a model identifies an intent from the given trajectory, we expect it to predict the most likely one, as reflected in the dataset distribution.
4. Evaluation Methodology
@@
This section outlines our evaluation methodology. Given an input UI trajectory and a corresponding gold task description, we assess whether a predicted task description matches the gold reference. We start with necessary definitions (Section 4.1), followed by a human and automatic evaluation protocols (Sections 4.2 and 4.3).
4.1. Definitions
@@
Inspired by the taxonomy proposed by Zhou et al. (2023), we differentiate between fulfillment for information-seeking intents and transactional intents (e.g. purchasing an item or changing settings). For the latter, fulfillment is achieved upon successfully completing the specific requirement outlined in the task. For information-seeking intents, fulfillment is achieved when the trajectory provides the necessary information sought in the user intent. We note that a fulfilling trajectory may provide some extra information beyond the intent, if such additional information is inherently bundled in the UI environment together with the sought information.
@@
We next aim to specify a matching criterion between two task descriptions, specifically a predicted one and the corresponding gold-reference. Given two task descriptions A and B and a UI environment, we say that A satisfies B in that environment if every reasonable trajectory that fulfills A would also fulfill B. In essence, this means that completing task A necessarily results in completing task B, making B a more general task than A in that UI environment. For instance, the task “Purchase the earliest train ticket to Edinburgh” satisfies the task “Purchase a train ticket to Edinburgh” but not vice versa.
@@
4.2. Human Evaluation Protocol
As with many text generation tasks, human evaluation is essential due to the limited reliability of automatic evaluation metrics. In our case, the annotator observes the gold and predicted task descriptions, with the corresponding trajectory, and assesses (1) whether the trajectory fulfills the predicted task, and (2) whether each of the predicted and gold task descriptions satisfies the other. Additionally, annotators validated the given data instance, excluding from the evaluation noisy instances in the original datasets (details in Appendix B).
To assess the quality of our proposed metric, we randomly sampled 50 instances from each of our datasets (see § 5.1) and generated task descriptions using our two baseline models (see § 5.2). We then measured the pairwise inter-annotator agreement among three of the authors, resulting in an average Cohen’s Kappa of 0.79 and 0.77 for fulfillment and satisfaction judgments, respectively, in the web dataset, and 0.91 and 0.86 in the Android dataset. These agreement levels are considered high according to Kappa values, which justified our decision to manually evaluate each baseline model in our experiments using a single annotator.
4.3. Automatic Evaluation Metric
We propose utilizing a Large Multimodal Model (LMM) as an automatic evaluator for the satisfaction criteria. Recent advancements in LMMs, such as those demonstrated in (He et al., 2024) and (Pan et al., 2024), have shown promising results in employing GPT-4 (Achiam et al., 2023) to assess task completion by autonomous agents. Building on this, we leverage the latest GPT-4o model as the automatic evaluator, to determine whether two task descriptions are mutually satisfied in the context of the trajectory (prompt details are in Appendix D). Measuring agreement with our human evaluation yielded a Kappa value of 0.48 (moderate agreement), suggesting a potential utility of model-based automatic evaluation for development cycles while highlighting the need for manual evaluation.
Table 1. Manual evaluation of human-generated task descriptions, at the different levels of match criteria (determined by the satisfaction relation) between annotators (Hi) and gold task descriptions (G), as well as among annotators. (H) represents average scores.
Dataset Non Match Partial Match Match (↑)
@@ 4.3. Automatic Evaluation Metric
AitZ H vs. G 0.02 0.22 0.76
   H1 vs. H2 0.02 0.04 0.94
5. Datasets and Baseline Models
Table 2. Manual and automatic evaluation scores of fulfillment relation and the different levels of match criteria (determined by the satisfaction relation), between model predictions and gold task descriptions.
@@ 5. Datasets and Baseline Models
Dataset Model Fulfillment Non Match Partial Match Match (↑) Non Match Partial Match Match (↑)
@@
AitZ GPT-4 0.78 0.22 0.19 0.59 0.32 0.34 0.34
Gemini-1.5 0.88 0.17 0.26 0.57 0.37 0.27 0.36
5.1. Evaluated Datasets
A suitable dataset for our task should contain natural language descriptions of user intents, along with traces of UI sessions that fulfill them. Aiming to satisfy this requirement, leveraging datasets that were created originally for UI automation is a natural choice. In these datasets, each instance was created by first specifying a particular user intent, followed by performing and recording a UI session that fulfills the given intent. Thus, a suitable dataset for our task of intent recognition is obtained from a UI automation dataset by swapping the input and output roles.
To that end, we explored several existing UI automation datasets but found that many of them do not fully align with the requirements of generic real-world intent recognition. For example, WebGUM operates in a simplified GUI environment while WebShop (Furuta et al., 2023; Yao et al., 2022) is limited to the shopping domain. RUSS and WebArena (Xu et al., 2021; Zhou et al., 2023), while providing user intents or instructions, do not possess the needed UI traces. PixelHelp and OmniACT (Li et al., 2020; Kapoor et al., 2025) provide only UI-level procedural instructions rather than natural high-level intents which are independent of a particular UI environment (asking to click a button, filling a text field etc.). Lastly, RICO (Deka et al., 2017) lacks annotations for user intents.
@@ 5.1. Evaluated Datasets
Considering these limitations, we turned to datasets that do satisfy our task’s requirements, focusing on two UI environments: web and Android. For the web environment, we utilize the Mind2Web dataset (Deng et al., 2023), the most widely used benchmark for autonomous web agents. The dataset was created by curating diverse tasks across popular websites, with annotators performing a series of actions to complete the goal. For Android, we used the prominent Android in the Wild (AitW) (Rawles et al., 2023) dataset, while focusing on its quality-filtered subset Android in the Zoo (AitZ) (Zhang et al., 2024) (details in Appendix A).
5.2. Models
@@
Given the multimodal nature of UI trajectories, our models must be adept at interpreting both text and images. We selected two state-of-the-art LMMs, Gemini 1.5 Pro (Reid et al., 2024) and GPT-4. These models are at the forefront of handling combined text and image inputs and offer a sufficiently large context window for our experiments.
@@
In our experiments, the model was guided through a Chain-of-Thought (Wei et al., 2022) process, to analyze the trajectory in a step-by-step manner (prompt in Appendix D).
@@
In web, inspired by SeeAct (Zheng et al., 2024) we drew a red bounding box around the element that the user interacted with to guide the model’s attention, as well as to break ambiguity where textual action descriptions were not sufficient. To further focus models, we also truncated lengthy web images based on the bounding box position. (details in Appendix B). For Android, no special modifications needed, actions are overlaid on the screenshots, and the screenshots are naturally smaller.
6. Experiments
6.1. Human Performance Evaluation